# Intertwined order of generalized global symmetries

### v1

@@ 2.1 Model
To study the interplay of zero-form and one-form global symmetries, we should use a model that couples these symmetries together. We thus consider a model of a $\mathbb{Z}_{N}$ lattice gauge theory coupled to $\mathbb{Z}_{N}$ matter on a three-dimensional cubic lattice in Euclidean spacetime [^3] See Refs. [45, 46, 47] for a review of lattice gauge theory.. However, the matter is not minimally coupled to the gauge field because this type of coupling removes the global symmetries we are interested in examining. Instead, we place the gauge fields $a_{\mu}$ on links of the lattice, matter fields $\varphi$ on sites of the dual lattice (i.e., centers of cubes on the direct lattice), and couple the matter to the gauge fields using a lattice analogue of a topological term. This choice of model was largely inspired by previous work of the authors [35] on the Cardy-Rabinovici model [40, 41, 48, 49, 50], a (3+1)d $\mathbb{Z}_{N}$ lattice gauge theory with an analogue of a theta term. Here, we consider a partition function in the Villain form [51], given by
@@
where $\Delta_{\mu}\varphi(R)\equiv\varphi(R+e_{\mu})-\varphi(R)$ denotes the lattice difference along the link between site $R$ and site $R+e_{\mu}$. The gauge field $a_{\mu}\in\mathbb{R}$ and electric particle worldlines $n_{\mu}\in\mathbb{Z}$ are on links $\ell$ between sites $r$, and the Dirac strings $s_{\mu\nu}\in\mathbb{Z}$ of the monopoles live on plaquettes $P$. Meanwhile, $\varphi\in\mathbb{R}$ and $n\in\mathbb{Z}$ live on sites $R$ of the dual lattice, and the $s_{\mu}\in\mathbb{Z}$ are on links $\tilde{\ell}$ of the dual lattice. The last term, a (2+1)d analogue of the theta term in (3+1)d, is an interaction between each plaquette and the link on the dual lattice that intersects it (see Figure 1).
@@ 2.2 Global symmetries
As mentioned previously, our model is equipped with two global symmetries, which we now discuss more formally. There is an ordinary (i.e., zero-form) $\mathbb{Z}_{N}$ global symmetry under which $\varphi$ transforms as
@@
$V(R)=e^{i\varphi(R)}$ (2.8)
on a dual lattice site $R$. Under the symmetry transformation, Eq. (2.7), this operator transforms as
@@
which is a product over all links $\ell$ which define the noncontractible loop $\Gamma$. The one-form global symmetry, Eq. (2.10), transforms the Wilson loop to
@@
$A_{\mu}\to A_{\mu}+\Delta_{\mu}\chi+N\mathcal{N}_{\mu},\qquad\varphi\to\varphi+\frac{2\pi}{N}\xi,\qquad s_{\mu}\to s_{\mu}-\mathcal{N}_{\mu},$ (2.15)
@@ 2.3 Duality
Before turning to the phase diagram of our lattice model, Eq. (2.1), it is useful to determine any non-perturbative information we can exploit to constrain the phase diagram. The model turns out to have two kinds of duality transformations. First, we note that in the absence of background fields, the partition function is periodic under shifts of $\Theta$ by integer multiples of $2\pi$. If we also introduce background gauge fields and perform the shift $\Theta\to\Theta+2\pi$, the action in Eq. (2.13) changes by
@@
Because the background fields are locally flat, Eq. (2.19) is a shift of $n$ and $n_{\mu}$ by integers, so such a transformation is well-defined. After the shift $\Theta\mapsto\Theta+2\pi$, the only term we cannot remove is
@@
Next, we turn to the other duality transformation, which is analogous to Kramers-Wannier duality. As demonstrated in Appendix C, by manipulating the partition function, Eq. (2.13), in ways similar to Refs. [7, 53], we can find a dual theory with action,
@@
Secondly, a more subtle difference is that the flat background gauge fields in Eq. (2.13) are now dynamical in Eq. (2.21), so the lattice model is dual to a gauged version of itself. This duality is a generalization of the duality relating a $\mathbb{Z}_{N}$ clock model and a $\mathbb{Z}_{N}$ gauge theory in (2+1)d.
@@
where $a_{\mu}$ and $b_{\mu\nu}$ are dynamical $\mathbb{Z}_{N}$ gauge fields while $A_{\mu}$ and $B_{\mu\nu}$ are background $\mathbb{Z}_{N}$ gauge fields. We will use both Eqs. (2.26) and (2.27) interchangeably. The mathematical structure of the $\mathcal{S}$ and $\mathcal{T}$ transformations is reminiscent of that in the fractional quantum Hall effect and (2+1)d conformal field theories [54, 55]. A similar structure also arises in (3+1)d lattice gauge theory with a theta term [41].
As we demonstrate in Appendix C, we can deduce how the $\mathcal{S}$ and $\mathcal{T}$ transformations map between different operators, allowing us determine what objects condense in a given phase. The local operators are labeled $(q_{s},q_{m})$ by a spin charge $q_{s}$ and a monopole charge $q_{m}$, and the loop operators are labeled $(q_{e},q_{v})$ by an electric charge $q_{e}$ and a vorticity $q_{v}$. Because the dynamical spins and electric charges are multiples of $N$, we note that a local operator that can condense must have $q_{s}\in N\mathbb{Z}$, and a loop operator that condense must have an electric charge $q_{e}\in N\mathbb{Z}$. The $\mathcal{S}$ and $\mathcal{T}$ transformations act on these objects as
@@ 2.4 Phases and phase diagram
Figure 2: The phase diagram at $\Theta=0$ for fixed $\kappa=e/g$ as $\mathrm{Im}(\tau)=2\pi/Nge$ is varied. In this limit, the lattice model reduces to a decoupled $\mathbb{Z}_{N}$ clock model and $\mathbb{Z}_{N}$ gauge theory. At large $e$ and $g$, there is a trivial phase in which the clock model is disordered and the gauge theory is confining, fully preserving the global symmetry $G=\mathbb{Z}_{N}^{(0)}\times\mathbb{Z}_{N}^{(1)}$. At small $e$ and $g$, the clock model orders, and the gauge theory is deconfined (and topologically ordered), breaking the full symmetry group $G$. Depending on the value of $\kappa=e/g$, there may be an intermediate phase in which one of $\mathbb{Z}_{N}^{(0)}$ or $\mathbb{Z}_{N}^{(1)}$ is broken but not the other.

### v2

@@ 2.1 Model
To study the interplay of zero-form and one-form global symmetries, we should use a model that couples these symmetries together. We thus consider a model of a $\mathbb{Z}_{N}$ lattice gauge theory coupled to $\mathbb{Z}_{N}$ matter on a three-dimensional cubic lattice in Euclidean spacetime [^4] See Refs. [49, 50, 51] for a review of lattice gauge theory.. However, the matter is not minimally coupled to the gauge field as in Refs. [52, 53] because this type of coupling removes the global symmetries we are interested in examining. Instead, we place the gauge fields $a_{\mu}$ on links of the lattice, matter fields $\varphi$ on sites of the dual lattice (i.e., centers of cubes on the direct lattice), and couple the matter to the gauge fields using a lattice analogue of a topological term. This choice of model was largely inspired by previous work of the authors [37] on the Cardy-Rabinovici model [43, 44, 54, 55, 56], a (3+1)d $\mathbb{Z}_{N}$ lattice gauge theory with an analogue of a theta term. The partition function for the lattice model we consider, first introduced by Shapere and Wilczek [38], is given by
@@
where $\Delta_{\mu}\varphi(R)\equiv\varphi(R+e_{\mu})-\varphi(R)$ denotes the lattice difference along the link between site $R$ and site $R+e_{\mu}$. For technical reasons, we work in the Villain form [57, 58] of the theory. The gauge field $a_{\mu}\in\mathbb{R}$ and electric particle worldlines $n_{\mu}\in\mathbb{Z}$ are on links $\ell$ between sites $r$, and the Dirac strings $s_{\mu\nu}\in\mathbb{Z}$ of the monopoles live on plaquettes $P$. Meanwhile, $\varphi\in\mathbb{R}$ and $n\in\mathbb{Z}$ live on sites $R$ of the dual lattice, and the $s_{\mu}\in\mathbb{Z}$ are on links $\tilde{\ell}$ of the dual lattice. The last term, a (2+1)d analogue of the theta term in (3+1)d, is an interaction between each plaquette and the link on the dual lattice that intersects it (see Figure 1).
@@ 2.2 Global symmetries
As mentioned previously, the lattice model is equipped with two global symmetries, which we now discuss more formally. There is an ordinary (i.e., zero-form) $\mathbb{Z}_{N}$ global symmetry under which $\varphi$ transforms as
@@
$V(R)=e^{i\varphi(R)},$ (2.8)
which is defined on a dual lattice site $R$. Under the symmetry transformation, Eq. (2.7), this operator transforms as
@@
which is a product over all links $\ell$ that define the noncontractible loop $\Gamma$. The one-form global symmetry, Eq. (2.10), transforms the Wilson loop to
@@
$A_{\mu}\to A_{\mu}+\Delta_{\mu}\chi+N\mathcal{N}_{\mu},\qquad\varphi\to\varphi+\frac{2\pi}{N}\chi,\qquad s_{\mu}\to s_{\mu}-\mathcal{N}_{\mu},$ (2.15)
@@ 2.3 Duality
Before turning to the phase diagram of our lattice model, Eq. (2.1), it is useful to determine any non-perturbative information we can exploit to constrain the phase diagram. The model turns out to have two kinds of duality transformations [38], which we make more precise in this section. First, we note that in the absence of background fields, the partition function is periodic under shifts of $\Theta$ by integer multiples of $2\pi$. If we also introduce background gauge fields and perform the shift $\Theta\to\Theta+2\pi$, the action in Eq. (2.13) changes by
@@
Because the background fields are locally flat, Eq. (2.19) is a shift of $n$ and $n_{\mu}$ by integers, so this transformation is well-defined. After the shift $\Theta\mapsto\Theta+2\pi$, the only term we cannot remove is
@@
Next, we turn to the other duality transformation, which is analogous to Kramers-Wannier duality. As demonstrated in Appendix B, by manipulating the partition function, Eq. (2.13), in ways similar to Refs. [7, 60], we can find a dual theory with action,
@@
Secondly, a more subtle difference is that the locally flat background gauge fields in Eq. (2.13) are now dynamical in Eq. (2.21), so the lattice model is dual to a gauged version of itself. This duality is a generalization of the duality relating a $\mathbb{Z}_{N}$ clock model and a $\mathbb{Z}_{N}$ gauge theory in (2+1)d.
@@
where $a_{\mu}$ and $b_{\mu\nu}$ are dynamical $\mathbb{Z}_{N}$ gauge fields while $A_{\mu}$ and $B_{\mu\nu}$ are background $\mathbb{Z}_{N}$ gauge fields. We will use both Eqs. (2.26) and (2.27) interchangeably. The mathematical structure of the $\mathcal{S}$ and $\mathcal{T}$ transformations is reminiscent of that in the fractional quantum Hall effect [61] and (2+1)d conformal field theories [62]. A similar structure also arises in (3+1)d lattice gauge theory with a theta term [44].
As we demonstrate in Appendix C, we can deduce how the $\mathcal{S}$ and $\mathcal{T}$ transformations map between different operators, allowing us determine what objects condense in a given phase. The local operators are labeled $(q_{s},q_{m})$ by a spin charge $q_{s}$ and a monopole charge $q_{m}$, and the loop operators are labeled $(q_{e},q_{v})$ by an electric charge $q_{e}$ and a vorticity $q_{v}$. Because the dynamical spins and electric charges are multiples of $N$, we note that a local operator that can condense must have $q_{s}\in N\mathbb{Z}$, and a loop operator that can condense must have an electric charge $q_{e}\in N\mathbb{Z}$. The $\mathcal{S}$ and $\mathcal{T}$ transformations act on these objects as
@@ 2.4 Phases and phase diagram
Figure 2: The phase diagram at $\Theta=0$ for fixed $\kappa=e/g$ as $\mathrm{Im}(\tau)=2\pi/Nge$ is varied. In this limit, the lattice model reduces to a decoupled $\mathbb{Z}_{N}$ clock model and $\mathbb{Z}_{N}$ gauge theory. At large $e$ and $g$, there is a trivial phase in which the clock model is disordered and the gauge theory is confining, fully preserving the global symmetry $G=\mathbb{Z}_{N}^{(0)}\times\mathbb{Z}_{N}^{(1)}$. At small $e$ and $g$, the clock model orders, and the gauge theory is deconfined (and topologically ordered), breaking the full symmetry group $G$. Depending on the value of $\kappa=e/g$, there may be an intermediate phase in which either $\mathbb{Z}_{N}^{(0)}$ or $\mathbb{Z}_{N}^{(1)}$ is broken but not the other.